A certain circuit has equivalent impedance Z = 8+j4 Ω, and an applied voltage 42.5 cos( 1000t + 30°)

The power factor is:

 

 Select one:

 a 0.45

 b 0.8

 c 0.89

 d 0.32

 e 0.6


1
Expert's answer
2021-06-02T07:43:12-0400

The impedance is


"Z=R+jX,\\
|Z|=\sqrt{R^2+X^2}."

The power factor is


"\cos\phi=\frac{R}{|Z|},\\\space\\
\cos\phi=\frac{8}{\sqrt{8^2+4^2}}=0.89."